### v2

# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

Abstract
We present ACDiT, a novel Autoregressive blockwise Conditional Diffusion Transformer, that innovatively combines autoregressive and diffusion paradigms for modeling continuous visual information. By introducing a block-wise autoregressive unit, ACDiT offers a flexible interpolation between token-wise autoregression and full-sequence diffusion, bypassing the limitations of discrete tokenization. The generation of each block is formulated as a conditional diffusion process, conditioned on prior blocks. ACDiT is easy to implement, as simple as creating a Skip-Causal Attention Mask (SCAM) on standard diffusion transformer during training. During inference, the process iterates between diffusion denoising and autoregressive decoding that can make full use of KV-Cache. We show that ACDiT performs best among all autoregressive baselines under similar model scales on image and video generation tasks. We also demonstrate that benefiting from autoregressive modeling, pretrained ACDiT can be transferred in visual understanding tasks despite being trained with the diffusion objective. The analysis of the trade-off between autoregressive modeling and diffusion demonstrates the potential of ACDiT to be used in long-horizon visual generation tasks. We hope that ACDiT offers a novel perspective on visual autoregressive generation and unlocks new avenues for unified models.
1 Introduction
The concept of predicting the future has been a fundamental principle of Artificial Intelligence, ranging from the “next token prediction” objective of autoregressive language modeling [55, 6, 1] to predicting the next action in Reinforcement Learning [48, 62, 12, 57]. The success of Large Language Models (LLMs) [70, 71, 6, 1, 34, 2, 80] exemplify the power of such philosophy, demonstrating that complex abilities can arise from this simple objective.
Figure 1: The generation process of ACDiT. Pixels in each block are denoised simultaneously conditioned on clean context which are autoregressively generated. The block can be flexibly defined, such as part of pixels in images or several frames in videos.
However, the multimodal aspects of information beyond language have yet to fully capitalize on this paradigm. In the realm of visual generation, diffusion models [31, 63, 14, 64] have demonstrated superior generative capabilities, producing creative outputs that are virtually indistinguishable from human-generated content, as evidenced by innovations like Sora [5] and Stable Diffusion [58, 53, 20]. Notwithstanding their remarkable achievements, these models operate in a non-autoregressive manner. Diffusion models receive corrupted target sequences with complete length and reconstruct the intended output through in-place iterative refinement. The term “in-place” underscores the distinct nature of this approach compared to the autoregressive model, which provides subsequent prediction, thereby extending the sequence and progressing towards the future. Such variation makes the model struggle to learn temporal correlation in vision data and poses difficulties for developing integrated frameworks capable of seamlessly bridging the vision foundation model with unified multimodality modeling [17, 68, 87, 75, 79, 78] and world model [81, 19, 7, 88, 36, 77].
Existing works attempt to build the visual autoregressive model by converting visual generation tasks into discrete token prediction tasks with vector quantization techniques [21, 68, 75, 69] and training with a next-token prediction objective. However, approaching the continuous distribution requires huge vocabulary sizes and a high utilization rate [85, 76], which is a complex objective.
In this paper, we propose ACDiT, an Autoregressive blockwise Conditional Diffusion Transformer that fuses the diffusion process with the autoregressive paradigm. Instead of relying on discrete tokens, we demonstrate that visual autoregressive generation can also be achieved using continuous visual features. At a high level, we extend the autoregressive units from the individual discrete token to blocks, where each block consists of visual patches of flexible size. The generation of each block can be formulated as a conditional diffusion process based on the previous block. As Fig. 1 shows, for image generation, a block represents a small region of the image, whereas in video generation, a block can correspond to a single frame or multiple frames.
ACDiT is easy to implement, as simple as adding a Skip-Causal Attention Mask to the current diffusion transformer [52] architecture during training. The inference process is formatted as an iteration between the conditional diffusion denoising process within a block, conditioned on the complete clean context, and autoregressive generation of a new block appended as the new context. In this way, KV-Cache can be used for faster inference. In general, ACDiT offers the following inherent advantages: (i) ACDiT simultaneously learns the causal interdependence across blocks with autoregressive modeling and the non-causal dependence within blocks with diffusion modeling. (ii) ACDiT is endowed with clean continuous visual input, which eliminates the requirement for vector quantization and improves the transfering to visual understanding tasks after being trained with generative objectives. (iii) ACDiT makes full use of KV-Cache for flexible autoregressive generation in any length and can potentially benefit from other latest long-context techniques in text for long video generation.
We validate ACDiT in both image and video generation tasks on ImageNet [60] and UCF101 [65], respectively. Experimental results demonstrate that ACDiT outperforms all autoregressive baselines of the same model scale and achieves visual quality comparable to full-sequence diffusion models while exhibiting significantly higher inference speed when extended to long sequences. We hope ACDiT presents a fresh perspective on visual autoregressive generation and paves the way for exploring unified multimodal and world models. The codes and models are available at https://github.com/thunlp/ACDiT.
2 Related Work
Diffusion Models. The field of image generation has witnessed remarkable advancements with the introduction of diffusion models [31, 63, 14, 49]. U-Net [59] is the early mainstream choice of network architecture [64, 49, 58, 53]. Following that, Transformer [73] is applied to diffusion models for image generation, with groundbreaking work such as DiT [52] and U-ViT [3] marking significant milestones. A series work, including PixArt-$\alpha,\delta,\Sigma$[9, 11, 10], demonstrate the capability of DiT on text-to-image tasks. Additionally, several studies have applied DiT to video generation, such as Lumiere [4] and Movie Gen [54].
Autoregressive Generation on Discrete Tokens. The iGPT [13] first proposes autoregressively generating raw image pixels as a raster-scan sequence. Most visual autoregressive models [56, 75, 83, 15, 16] follow VQGAN [21], which pioneers this direction by training an autoregressive transformer on discrete tokens produced by VQVAE [72]. LlamaGen [67] enhances the image tokenizer and scales up the autoregressive transformers building on the latest Llama architecture [70]. Inspired by RQ-Transformer [38], VAR [69] proposes the next-scale prediction and obtain good improvement. Some subsequent works [69, 43, 82] extend next-scale prediction to text-to-image generation.
Autoregressive Generation on Continous Features. Some recent works make some early exploration on achieving visual autoregressive generation without discrete tokens and combining the advantage of diffusion and autoregressive. Diffusion Forcing [8] trains a causal autoregressive model to generate blocks without fully diffusing past ones and implement it on small RNN. MAR [41] proposes the diffusion loss to learn the autoregressive conditional distribution on the head of the main Transformer with a small MLP network. Transfusion [87] and Monoformer [86] conduct joint training with language modeling loss and diffusion loss in a single transformer, while within image generation, their design is identical to the standard diffusion model. Dart [25] proposes to autoregressively denoised the complete image within in a Transformer. Different from these work, ACDiT redefine the autoregressive unit and generates each block based on clear past, providing a flexible framework for both image and video generation.
(a) Skip-Causal Attention Mask.
(b) Inference process of ACDiT.
(c) 3D view of ACDiT.
Figure 2: (a): for each noised block $n_{i}$, it can only attend previous clean latent $c_{0},c_{1},\dots,c_{i-1}$ and itself. Each clean block $c_{i}$ only attends the previous clean latent block. (b): ACDiT can effectively utilize the KV-Cache for autoregressive inference. (c): the 3D view of ACDiT, where $B,L,T$ denotes the block size, number of blocks, and denoising time steps, respectively. The darker color denotes higher noises.
3 Prerequiste
3.1 Autoregressive Modeling
Autoregression asserts that the value at each timestep is contingent upon its preceding values. This principle is exemplified in autoregressive language models, which iteratively predict the probability distribution of subsequent tokens. Given a sequence of tokens ($x_{1},x_{2},\dots,x_{n}$), a salient characteristic of autoregression is that the prediction of $x_{i}$ is only dependent on its prefix ($x_{1},x_{2},\dots,x_{i-1}$). Upon determining $x_{i}$, it is concatenated with the preceding sequence, thereby forming the conditioning context ($x_{1},x_{2},\dots,x_{i}$) for predicting $x_{i+1}$. Therefore, the likelihood of sequence can be factorized as:
$p(x_{1},x_{2},\dots,x_{n})=\prod_{i=1}^{n}p(x_{i+1}|x_{1},x_{2},\dots,x_{i}).$ (1)
Thanks to the flexibility of self-attention in Transformers, autoregressive models can be effectively implemented by adding a causal mask in Transformer attention block [73].
3.2 Diffusion
Diffusion models, in contrast, conceptualize a noise-infusion and denoising process, which is defined by gradually adding noise to the initial data $n_{0}$ and training the model to learn the inverse mapping. Formally, the noised data $n^{(t)}$ at each step $t$ is sampled by $q(n^{(t)}|n^{(t-1)})=\mathcal{N}(x_{t};\sqrt{\alpha^{(t)}}n^{(0)},(1-\alpha^{(t)})\mathbf{I})$ [^1] For clarification, we use subscript t to denote timesteps in autoregressive models and superscript (t) to denote the timesteps in diffusion models., which is equivalent to add a Guassain noise to the previous samples: $n^{(t)}=\sqrt{\alpha^{(t)}}n^{(t-1)}+\sqrt{1-\alpha^{(t)}}\epsilon^{(t)},\epsilon^{(t)}\sim\mathcal{N}(0,\mathbf{I})$ while the diffusion model $p_{\theta}$ is trained to learn the reverse process $p_{\theta}(n^{(t-1)}|n^{(t)})=\mathcal{N}(\mu_{\theta}(n^{(t)}),\beta^{(t)}\mathbf{I})$. With the reparameterization trick, the network $\mu_{\theta}(n^{(t)})$ can be reparameterized as noise prediction network $\epsilon_{\theta}(n^{(t)})$ and the training objective can be simple as:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\sim U[0,1],\epsilon\sim\mathcal{N}(0,I)}||\epsilon_{\theta}(n^{(t)},t)-\epsilon^{(t)}||_{2}.$ (2)
During the inference phase, the denoising process is initialized with a random Gaussian noise sample $n^{(T)}$, followed by $T^{\prime}$ denoising steps, ultimately yielding a single deterministic samples $\tilde{n}^{(0)}$ from its underlying distribution. Typically, the denoising process in diffusion models operates “in-place”, meaning that each new denoising step directly replaces the previous step’s input. This differs from autoregressive modeling, where the value of a subsequent step is appended to the existing sequence.
3.3 Desiderata
A robust autoregressive diffusion method should integrate the strengths of both autoregressive modeling and diffusion. To achieve this synergy, we have identified three critical desiderata that the framework must meet:
  1. 1.
    The generation of future elements should be predicated on a precise representation of antecedent sequences. This is imperative because any ambiguity in the past inevitably complicates future predictions. This approach preserves the efficacy of autoregressive modeling and potentially facilitates the development of the world model. Furthermore, adherence to this principle enhances performance in discriminative tasks (e.g., visual understanding), as these tasks necessitate the input of all observable features into the model.
  2. 2.
    Both the autoregressive modeling and the denoising process should optimally utilize the entire parameter space of the neural network. In an elegant fusion of autoregressive models and diffusion, neither component should be relegated to an auxiliary role. Instead, they should function as integral, complementary elements of the system.
  3. 3.
    The denoising process should directly attend comprehensively to the entire sequence of past sequences. Failure to do so would necessitate that the denoise process’s condition encapsulates all preceding information, placing an unrealistic demand on the lossless compression capabilities of the feature space. On the contrary, a holistic input of all past information in each denoise step ensures a more effective processing of temporal dependencies.
Based on these desiderata, we analyze representative existing autoregressive diffusion methods. Diffusion-Forcing [8] proposes to use different-level random noise in the different positions of a sequence. Thus, in the inference process, denoising the subsequent positions from a clean past can be seen as a special case denoising different levels of noise. Their method does not meet the first desideratum. In the training process, the future is not predicted from the precise representation of the past. In MAR [41], the diffusion process is trained based on the latent in the last position, which does not satisfy the third desideratum. Moreover, the diffusion process sorely leverages the head part of the network, which conflicts with the second desideratum. In Transfusion [87] and Monoformer [86], despite that the diffusion utilizes the full network parameters for both autoregressive modeling and diffusion, it does not utilize the “predicting the future” objective within the visual information. Moreover, when trained with multiple visual input, such as multiple images, the latter image will attend to the noise version of the former image. To comprise this deficiency, they half the noise schedule, i.e., limiting maximum noise steps to 500 from 1000, in 20% image captioning pairs, which is not an optimal strategy.
Figure 3: Left: an illustrative figure of the ideal dependencies between the noise and clean information, which is adopted by ACDiT. Right: the dependencies of diffusion head in MAR.
4 ACDiT
4.1 Framework
To satisfy the desiderata discussed above, we propose a versatile framework for autoregressive diffusion called ACDiT. For generality, ACDiT runs block-wise autoregression instead of token-wise autoregression. We identify there are two kinds of blocks: the clean blocks $c_{i}$ and the noise blocks $n_{i}$, where $n_{i}$ is corrupted from $c_{i}$. ACDiT learns several conditional distribution $p(c_{i}|c_{<i})$ factorized with Eq. 1, where $p(c_{i}|c_{<i})$ is optimized by learning the conditional noise prediction network $\epsilon_{\theta}(n_{i}^{(t)};t,c_{<i})$ with Eq. 2. The final training objective is:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\sim U[0,1],\epsilon\sim\mathcal{N}(0,I)}\sum_{i=1}^{n}||\epsilon_{\theta}(n_{i}^{(t)};t,c_{<i})-\epsilon^{(t)}||_{2}.$ (3)
Given the aforementioned desiderata, we can conceptualize all $n_{i}$ and all $c_{i}$ as occupying separate positions, effectively transforming the dependency structure into an attention pattern between different positions. We designate this attention pattern as the Skip-Causal Attention Mask (SCAM), which is visually represented in Figure 2(a). The figure elucidates that $n_{i}$ attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself, while $c_{i}$ also attends to all preceding clean blocks $\{c_{j}|j=0,...,i-1\}$ and itself. In training, for both simplicity and efficiency, we can group the attention mechanisms as illustrated in the right matrix of Figure 2(a). Suppose the number of blocks is $N$, then the unmasked positions form two triangular matrices of side length $N-1$, complemented by a diagonal matrix of side length $N$.
During the inference phase, each autoregressive step executes a conditional diffusion process for $n_{i}$ based on $\{c_{j}|j=0,...,i-1\}$’s KV-Cache. Upon finishing denoising, it is appended to the clean sequence as $c_{i}$ followed by the maximal noise version of the next block $n_{i+1}$. The key-value tensor will be computed for these two blocks, and the key-value tensor of the clean block $c_{i}$ will be kept in KV-Cache. All noise-corrupted version of $n_{i}$ is disregarded. The process is visualized in Figure 2(b). A three-dimensional view of our method is presented in Figure 2(c). By interpolating between full-sequence diffusion and autoregressive paradigms, ACDiT enjoys flexibility and expressivity, enabling it to generate video of any length utilizing the latest long-context techniques developed for language models.
4.2 Positional Encoding
ACDiT is designed to be versatile, capable of handling one-, two-, three-, or even higher-dimensional data, including but not limited to text (1D), images (2D), and video (3D). For any given dimension of data, the position of that data is a critical attribute that must be made known to the model. This positional awareness enables the model to contextualize its current focus relative to historical data. In the domain of textual data, the Rotary Position Embedding (RoPE) [66] has gained widespread adoption as an effective relative positional encoding method. To address the challenges posed by multi-dimensional positional indices, we introduce RoPE-ND, a natural extension of RoPE.
For a token of a $D$ dimensional data, its positional index is $[m_{1},m_{2},...,m_{D}]$. Given query and key vectors in Transformer’s attention module, we partition the hidden dimension into $D$ segments. It is imperative that each segment’s hidden dimension be an even number. For each segment $j$, we apply a RoPE with a specific base $b_{j}$, as defined in the following equation:
[…]
In this formulation, each $\mathbf{R}^{d_{j}}_{\Theta_{j},m_{j}}$ represents a $d_{j}$-dimensional rotary matrix [^2] See a detailed explanation in the Equation 15 in  [66]. with rotation angles $\Theta_{j}=\{\theta_{ji}=b_{j}^{-2(i-1)/d_{j}},i\in[1,2,\cdots,\frac{d_{j}}{2}]\}$. The base $b_{j}$ is empirically determined as $100\lceil\frac{8L_{j}}{100\pi}\rceil$, where $L_{j}$ denotes the maximum position index in data dimension $j$. This formulation ensures that the highest wavelength of RoPE is approximately eight times the maximum position, thereby mitigating rapid decay in long-term dependencies.
It is worth noting that ACDiT inherently supports length extrapolation [66], although a comprehensive exploration of this capability falls beyond the scope of the present work.
4.3 Efficiency Analysis and Block Size Choice
In this section, we provide a brief analysis on the computational efficiency of ACDiT in terms of floating-point operations (FLOPS). We first assume that denoising each block requires the same number of time steps $T$ as the full sequence diffusion, despite that when the block size is small, it may require fewer denoising steps, thus making ACDiT potentially more efficient. Let $\theta$ denote the number of parameters in one layer of the transformer block, $h$ represents the hidden dimension and $n$ as the number of attention heads. The FLOPS is formed by a $O(L)$ component $2L\theta$, from the product between input and weight matrices, and a $O(L^{2})$ component from the pairwise computation between Q-K pairs in the attention dot product, which is $4(h+n)$ for each Q-K pair in the $L^{2}$ Q-K pairs. Consequently, given that $n\ll h$, the total FLOPS can be expressed as: $F=2L\theta+4(h+n)L^{2}\approx 2L\theta+4hL^{2}$.
When employing ACDiT with KV-Cache, the FLOPS consumed by FFN and attention parameters remain unchanged. However, the attention mechanism transforms into causal attention with block size $B$. In this scenario, the number of Q-K pairs to be calculated is given by:
$\sum_{i=1}^{L/B}iB^{2}=\frac{1}{2}(\frac{L}{B}+\frac{L^{2}}{B^{2}})B^{2}.$ (5)
Thus, the FLOPS saved can be expressed as $4h{L^{2}}(\frac{1}{2}(1-\frac{B}{L}))$. Compared to the total FLOPS, the saved percentage is
$\frac{4h{L^{2}}(\frac{1}{2}-\frac{B}{2L})}{4hL^{2}+2L\theta}=\frac{1-B/L}{2+m/k},$ (6)
where $m=\frac{\theta}{h^{2}}$ and $k=\frac{L}{h}$. This equation demonstrates that when the sequence length significantly exceeds the hidden size, transforming a full sequence diffusion into ACDiT can save up to 50% of FLOPS.
However, in practice, it is not always beneficial for small $B$. Setting $B$ to an excessively small value may not fully leverage the iterative modification inherent in the diffusion process, potentially compromising generation quality. Furthermore, given the parallel computing nature of computational kernels, a very small $B$ may not yield speed improvements, a phenomenon analogous to the rationale behind speculative decoding [39]. Conversely, setting $B$ to a very large value diminishes efficiency both in terms of attention calculation, as shown in Equation 6. It also fails to capitalize on the strengths of auto-regressive generation. Indeed, when $B$ is set equal to $L$, ACDiT reverts to the original DiT model. We analyze the influence of block size on performance and efficiency in Sec. 5.3.
Figure 4: The model architecture of ACDiT, mainly inherited from standard DiT architecture. Both the clean latent and noise latent are input while only noise tokens are scale shift with conditioning information. The SCAM attention mask pattern will be applied in the attention block.
Model #Layers Hidden Size MLP #Heads Params
ACDiT-B 12 768 3072 12 132M
ACDiT-L 24 1024 4096 16 460M
ACDiT-XL 28 1152 4608 18 677M
ACDiT-H 32 1280 5120 20 954M
Table 1: Configuration details of ACDiT.
[…]
ADM [14] (NeuIPS’19) Diff. - - 554M 10.94 101.0 0.69 0.63
LDM-4-G [58] (CVPR’22) Diff. Cont. - 400M 3.60 247.7 - -
DiT-L/2 [52] (ICCV’23) Diff. Cont. - 458M 5.02 167.2 0.75 0.57
DiT-XL/2 [52] (ICCV’23) Diff. Cont. - 675M 2.27 278.2 0.83 0.57
MaskGIT [76] (CVPR’22) Mask. Disc. - 227M 6.18 316.2 0.83 0.58
MAGE [40] (CVPR’23) Mask. Disc. - 230M 6.93 195.8 - -
VQGAN [21] (CVPR’21) AR Disc. 1.4B 15.78 78.3 - -
RQTran [38] (CVPR’22) AR Disc. 3.8B 7.55 134.0 - -
VAR-d16 [69] (NeurIPS’24) VAR Disc. 310M 3.30 274.4 0.84 0.51
VAR-d20 [69] (NeurIPS’24) VAR Disc. 600M 2.57 302.6 0.83 0.56
LlamaGen-L [67] (arxiv’24) AR Disc. 343M 3.07 256.1 0.83 0.52
LlamaGen-XL [67] (arxiv’24) AR Disc. 775M 2.62 244.1 0.80 0.57
LlamaGen-XXL [67] (arxiv’24) AR Disc. 1.4B 2.34 253.9 0.80 0.59
ImageFolder [42] (ICLR’25) AR Disc. 362M 2.60 295.0 0.75 0.63
LlamaGen+NPP [50] (arxiv’24) AR Cont. 1.4B 2.55 282.0 0.84 0.56
MAR-L [69] (NeurIPS’24) AR Cont. 479M 4.07 232.4 - -
MAR-L [69] (NeurIPS’24) MAR Cont. - 479M 1.78 296.0 0.81 0.60
MonoFormer [86] (arxiv’24) AR + Diff Cont. - 1.1B 2.57 272.6 0.84 0.56
Dart [26] (ICLR’25) AR + Diff Cont. 820M 3.82 263.8 - -
ACDiT-L AR+Diff Cont. 460M 2.53 262.9 0.82 0.55
ACDiT-XL AR+Diff Cont. 677M 2.45 267.4 0.82 0.57
ACDiT-H AR+Diff Cont. 954M 2.37 273.3 0.82 0.57
Table 2: Image generation results on ImageNet 256x256.
4.4 Model Architecture
As shown in Figure 4, ACDiT mainly inherits the main architecture of DiT. We replace the original bidirectional attention with the proposed SCAM attention pattern to process the clean and noise latent. Since we want to keep the architecture as simple and unified as possible, we use linear layers instead of convolution in the input layer and final layer. Besides, we replace the absolute position embedding and Layer Normalization with RoPE [66] and RMSNorm (Root Mean Square Layer Normalization) [70], respectively. We find that QK-norm is important to stabilizing the video generation training, thus we use QK-norms in all experiments. The additional conditional information timesteps and labels are injected into the model with adaLN-Zero only on the noise part. For both image and video generation, we follow DiT and leverage the pre-trained image VAE [35] from Stable Diffusion [58], whose downsample factor is 8. For image generation under B block size, we group square latent representation patches with $\sqrt{B}\times\sqrt{B}$ shape as a block.
5 Experiments
5.1 Experimental Setup
Dataset. For image generation tasks, we consider the ImageNet [60] dataset with 256$\times$256 resolution, which consists of around 1.28M images from 1K classes. For video generation, we consider the UCF-101 [65] dataset with 16 frames, where each frame is an image with 256$\times$256 resolution. UCF-101 contains 13320 videos from 101 classes.
Implementation details. In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. Therefore, for a $256\times 256\times 3$ image in $32\times 32\times 4$ latent shape, the total sequence length and autoregressive length are 1024 and 4, respectively. We explore 4 different model sizes, as shown in Table 1. ACDiT-B is used for design verification and analyze. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer [44] and WSD (Warmup Steady Decay) learning rate scheduling [33] with the peak learning rate 3e-4 and no weight decay. The learning rate begins to decay in the last 15% training iteration. Following the common training recipe of generative models, we keep an exponential moving average (EMA) of the ACDiT weights during training using a decay rate of 0.9999. We sample images with DPM-Solver [45] for 25 steps within each block and use classifier-free guidance [30] with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. For a $16\times 256\times 256\times 3$ video in $16\times 32\times 32\times 4$ latent shape, the sequence length of each frame is 256 and the total sequence length is 4096, with 4 frames grouped into one block. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Other training configs are the same as image training. All models are implemented with PyTorch [51]. Specifically, we use FlexAttention [^3] https://pytorch.org/blog/flexattention to implement the SCAM for both customization and efficiency. We use 256 and 96 NVIDIA H100 GPUs to train the image and video generation model, respectively.
Figure 5: FID and FVD curves of ACDiT-B over training steps with different sequence lengths and autoregressive lengths. PS means patch size.
Figure 6: The change of inference time in different autoregressive lengths under varied total sequence length.
(a) Scaling performance of ACDiT.
(b) Ablation study for ROPE-ND.
(c) FID curve of last 30% training.
Figure 7: (a): ACDiT shows scaling performance similar to DiT. (b): ROPE-ND has consistent improvement to the generation quality. (c): FID score sharply with the learning rate beginning to decay when using the WSD scheduler.
[…]
LVDM [28] (arxiv’22) Diff. 437M 372
Latte [47] (arxiv’24) Diff. 674M 478
Matten [23] (arxiv’24) Diff. 853M 211
VideoFusion [46] (CVPR’23) Diff. 510M 173
MMVG [22] (CVPR’23) Mask. 230M 328
MAGVIT [84] (CVPR’23) Mask. 306M 76
MAGVITv2 [85] (ICLR’24) Mask. 307M 58
TATS [24] (ECCV’22) AR 331M 332
CogVideo [32] (ICLR’23) AR 9.4B 626
MAGVIT-AR [84] (CVPR’23) AR 306M 265
MAGVITv2-AR [85] (ICLR’24) AR 307M 109
OmniTokenizer [74] (NeurIPS’24) AR 650M 191
ACDiT-XL AR+Diff. 677M 111
ACDiT-H AR+Diff. 954M 104
ACDiT-H-LT AR+Diff. 954M 90
Table 3: Video generation results on the UCF-101 dataset. ACDiT-XL-LT means ACDiT-XL trained for longer epoch.
Model Type Top-1 Acc
ViT-H[18] Supervised 83.1
MAGE[40] Masked. 84.3
MAE[27] Masked. 85.9
iGPT [13] Generative 72.6
DiT-XL [52] Generative 82.8
ACDiT-XL Generative 84.0
Table 4: Supervised fine-tuned Top-1 accuracy on Imagenet. We finetune ACDiT-XL and DiT-XL for 100 epoch.
Figure 8: Sample images from ACDiT-H on ImageNet 256$\times$256.
Figure 9: Sample videos from ACDiT-XL trained on UCF-101.
5.2 Main Results
Image Generation. We report the FID-50K [29], Inception Score [61], Precision and Recall [37] of ACDiT and baselines in Table 2. Compared with previous autoregressive models and masked generative models utilizing discrete tokens, such as VQGAN, VAR, LlamaGen, and MaskGIT, ACDiT consistently achieves superior performance with lower FID scores at comparable model scales. Notably, ACDiT-XL achieves 2.45 FID scores, outperforming both LlamaGen-XXL and VAR-d20 with similar parameters. Additionally, when compared to the MAR-L variant that does not recompute attention, ACDiT-L significantly improves performance across all metrics. Although MAR-L has lower FID than ACDiT, recomputing attention makes it hard to generalize longer sequence generation. Compared with other autoregressive diffusion methods, such as Monoformer [86] and Dart [25], ACDiT has a significantly superior performance. When compared with leading diffusion-based methods, ACDiT also demonstrates competitive performance. For instance, despite not employing full-sequence attention, ACDiT models achieve results close to DiT-XL. In general, these results highlight the distinct advantages of ACDiT over other baselines with the continuous latent representation and KV-Cache. Qualitative results are presented in Appendix.
Video Generation. Different from image generation, video inherently includes a temporal dimension, making it more well-suited to autoregressive modeling. The FVD metric on UCF-101 for class-conditional video generation is reported in Table 3. With hybrid AR+Diff architecture, ACDiT-H achieves much lower FVD than other diffusion-based and autoregressive methods, even outperforming MAGVIT-AR and MAGVITv2-AR, which utilizes a closed-source, specially designed video tokenizer. In contrast, ACDiT simplifies the process by directly using an open-sourced image VAE. Although MAGVIT with masked generative methods has a lower FVD than ACDiT, they rely on ”in-place” operation to generate a video similar to the diffusion model. This constraint limits their ability to generalize to generate longer video generation and build world models. Compared to image generation, ACDiT demonstrates a greater potential in the more complex domain of video generation, where longer visual sequences and temporal information are critical. Qualitative results of ACDiT-XL are presented in Fig. 9.
Image Representation. We assess the capability of ACDiT in image representation, which is essential for building a unified visual understanding and generation model. We finetune ACDiT-XL and DiT-XL on ImageNet using classification loss following the training setting in MAE [27] and report the Top-1 accuracy in Table 4. The accuracy of ACDiT surpasses that of ViT-H, iGPT and DiT-XL. This superior performance over DiT-XL highlights the benefit of incorporating clean latent inputs, which accelerates the model’s ability to learn better representations compared to using only noised latent inputs. Furthermore, ACDiT is on par with MAGE [40] in terms of Top-1 accuracy, while enjoying better generation capabilities.
5.3 Analysis
Trade off of block size. Fig. 5 illustrates the trend of trade-off under different sequence lengths and block sizes in image and video generation tasks on ACDiT-B. The FID curve indicates that for image generation, directly increasing the autoregressive length leads to a decline in image quality, since each patch receives less attention information on average. However, we can mitigate this decline by increasing the total sequence length, which means reducing the patch size. For video generation, ACDiT shows more advantages due to the inherent temporal dependence of videos. The FVD curve demonstrates that increasing autoregressive length has minimal effect on the video quality, even with slight improvement. As for efficiency, we test the sampling time for various sequence lengths with a batch size of 4 on an NVIDIA A100 GPU. Fig. 6 shows that as the sequence length increases, particularly beyond 16k, full-sequence attention (AR length of 1) becomes very time-consuming, necessitating the autoregressive generation.
Scaling Performance. We present the scaling performance of ACDiT in Fig. 7(a). For a fair comparison with DiT, we use the same batch size and learning rate as DiT in these training sessions. When increasing the model size, ACDiT shows consistent improvement in image quality across all autoregressive lengths, sharing a similar scaling trend with DiT. Notably, the improvement is more pronounced with longer autoregressive lengths. We hypothesize that this is due to reduced accumulation of errors when scaling the model size.
Ablation Study. We ablate the effectiveness of ROPE-ND positional embedding on image generation with patch size as 2. As shown in Fig 7(b), adding ROPE-ND results in consistent improvements in FID score.
Training dynamics of WSD scheduler. Unlike the constant learning rate used in DiT, we utilize the WSD learning rate scheduler [33]. WSD scheduler maintains a constant learning rate as the main stage of training, while one can diverge from the main branch at any time, potentially based on the compute budget, with a rapidly decaying learning rate. As Fig. 7(c) shows, the FID remains almost converged during the constant learning rate state, while sharply dropping after the learning decays, similar to the loss curve when using the WSD scheduler in LLM training. To the best of our knowledge, we are the first to validate the effectiveness of the WSD scheduler in visual generation.
6 Conclusion
In this paper, we propose ACDiT that interpolates the autoregressive modeling and diffusion transformers. With a simple but novel design of attention mask, ACDiT can achieve autoregressive generation on any length while maintaining a clear latent input potentially for adding a visual understanding task. By combining the advantages of both autoregressive and diffusion models, we demonstrate the performance and efficiency of ACDiT in both image and video generation tasks. We hope ACDiT can shed light on designing the new visual autoregressive model and building a unified multimodal model in the future.